A star rotates in a circular orbit about the center of its galaxy. The radius of the orbit is 7.9 x 1020 m, and the angular speed of the star is 1.8 x 10-15 rad/s. How long (in years) does it take for the star to make one revolution around the center?
Answer #1

angular speed = = 2/T

where T is time period for one revolution

Hence , T = 2/ = 2 / (1.810-15 ) = 3.491 1015 s

time period in years = ( 3.491 1015 ) / ( 3600 24 365 ) = 1.11 108 years